OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (1 point) 5x – y = –2 10x – 3y = –7

OpenStudy (anonymous):

From Eq 1 y=5x+2 Put it in Eq 2 10x-15x-6=-7 x=1/5

hero (hero):

5x – y = –2 10x – 3y = –7 1. Write the first equation in the following form: 5x + 2 = y 2. Substitute the expression 5x+2 in place of y in the second equation: 10x - 3(5x+2) = -7 3. Simplify the left side of that equation: 10x - 15x -6 = -7 -5x - 6 = -7 4. Continue Solving for x: -5x = -7+6 -5x = -1 x = 1/5 5. Substitute the value for x back into the first equation: 5(1/5) - y = -2 1 - y = -2 1 + 2 = y 3 = y 6. Check using the second equation: 10(1/5) - 3(3) = -7 2 -9 = -7 -7 = -7